OpenStudy (anonymous):

How would I go about solving this equation? (m+5)/b = (q+3) + (r - 1) The idea is that you can write a quotient like this: m/b = q + r And the question is to find the values possible for b and m, that if you increment m by 5 the quotient increments by 3 and the rest diminishes by 1

OpenStudy (anonymous):

\[\frac{ m+5 }{ b } = q + 3 + r - 1\]